OpenStudy (anonymous):

make it a factor/factoring 24xy-8xy^2=

OpenStudy (ash2326):

x(24y-8y^2)=0 x(y(24-8y)=0 x(y(3-y)=0 either x=0 or (24-8y^2)=0 solving the second one we get y=0 and y=3 so solution is x=0 or y=0,3

OpenStudy (pokemon23):

hey ash can yo come to my question